# IPPP/24/10

### v1

@@ 2.2 Creating non-Gaussian operations
The circuit parameters are then trained by a classical machine learning routine to fit the output of the circuit to a target state vector. At each step of the training, the loss is calculated as
@@ 3 Implementing time-evolution of wavefunctions
Having collected the required ingredients, we now present a Continuous-Variable Quantum Computing (CVQC) framework for simulating the time-evolution of quantum states under the influence of an arbitrary potential. Section 3.1 describes the principle whereby an arbitrary state can be made to evolve under the influence of an arbitrary Hamiltonian, by using an ancilla qumode initialised with a prescribed non-Gaussian state called the evolver-state. Then Section 3.2 discusses how the evolver state may itself be constructed using the measurement based approach described in Section 2.2.
@@ 3.1 Schrödinger evolving wavefunctions with arbitrary Hamiltonians
The goal is to be able to evolve any arbitrary in-state $|\psi_{\rm in}\rangle$ under the influence of the Hamiltonian, ${\mathcal{H}}$, for which one must devise a photonic circuit that will implement the Schrödinger evolution
@@
The techniques that will be developed here to do this are applicable to any potential, $V$, but in order to have a specific system in mind it is useful to focus on two specific cases. The first is the system with the potential
@@
The depth of the true vacuum of this potential is $V_{\rm min}\approx\varepsilon$, at leading order. The potential is shown in Figure 3(a). The interesting aspect of this potential is that the simple harmonic oscillator Gaussian ground state at $x=0$ is an approximate energy eigenstate which will decay by tunnelling through the barrier, and therefore the behaviour that will be recovered is characteristically “quantum”.
Figure 4: The potentials of focus in this study. In both cases the wave-function is initialised as a Gaussian state centred on the origin,.
@@
This potential is of interest because its expansion around $x=1$ is $V(x)=(x-1)^{2}/2+\ldots$, so to quadratic order it is the SHO potential. Therefore deviation from SHO behaviour is directly attributable to the higher order terms, and moreover this deviation comes from a potential that is not polynomial.
To implement the Schrödinger evolution of the wavefunction in Equation (16), there are two approximations that will be made. The first approximation is to work within the Fock truncation alluded to earlier. That is
@@
where $A_{n}$ is the coefficient of the Fock state $|n\rangle$, and where $n_{\textrm{max}}$ is as before the Fock state truncation. Thus in principle it is possible to determine the evolution in terms of a matrix acting on the Fock state coefficients, $A_{n}$, by expanding $\hat{x}$ and $\hat{p}$ in terms of the creation and annihilation operators as in Equation (3). Although such Fock truncation is not strictly speaking required for a particular photonic circuit to work, it is required to determine the parameters of the circuit itself, as will become clear.
@@
where the dots denote higher order commutators. Assuming that $[\hat{p}^{2},H_{1}(\hat{x})]\sim 1$, for such a trotterised evolution one therefore finds that an error accumulates in the exponent of order $N\delta t^{2}$ and hence one requires $\delta t\ll 1/t$ in natural units for the evolution to be accurate. Note that the Trotterisation error is retained as a product of unitary operators, so that it is also unitary***It could in principle be improved by subtracting the leading $\delta t^{2}$ error with more compound Trotter steps, however this will not be necessary..
@@
According to Equations (7) and (13), performing the various manipulations corresponding to these gates and then performing the homodyne measurement $y=0$ with the choice of parameters in Equation (27) yields
@@
All that remains is to choose the optimal form of the resource function $\left\langle y|\phi_{0}\right\rangle$ in order to maximally suppress the effect of the noise function. There are two possibilities: either one can choose a small value of $q$ which ‘freezes’ the value of the function $\left\langle xq|\phi_{0}\right\rangle$, and/or choose a flat resource function. It is convenient to adopt the top-hat function as the idealised resource function:
@@
$|\psi\mbox{\small$(\delta t)$}\rangle~{}=~{}\frac{\Theta(L/2-q|\hat{x}|)}{L}~{}e^{-i{\cal H}(\hat{x})\,\delta t}|\psi_{\rm in}\rangle~{},$ (30)
where $\Theta$ is the Heaviside function. Note that smaller values of $q$ allow larger domains in $x$ of valid evolution.
@@ 3.2 Learning the evolver-state
To fully implement the procedure outlined in Section 3.1 on a photonic device, the non-Gaussian evolver-state, $\left|\phi\right\rangle$, must be prepared. Here, the measurement-based approach from Section 2.2 will be used, in which the circuit parameters of a circuit such as that in Figure 3 are trained against the target state from Equation (32). To achieve a good fit to the target, a circuit constructed from three-qumodes and 10 iterations of the layer method will be used. At each layer, the displaced-squeezed states pass through an interferometer which entangles the system. Following in the rectangular architecture from Reference Clements2016 only three beamsplitters are required in each layer for the three-qumode case. Each gate operation has two parameters, thus the full circuit has 180 trainable variables. The values that the PNR measurements are post-selected on have not been included as trainable parameters and have instead been chosen to be $m_{0}=m_{1}=5$. A schematic of the circuit diagram is shown in Figure 6.
@@
4 Results: QM on photonics versus numerically evaluated QM
@@
Figure 8: The time-evolution of a quantum system with the asymmetric potential of Equation (17) with (a) $\varepsilon=0.1$ and (b) $\varepsilon=0.5$, generated by the photonic quantum simulator with a Fock truncation of 60 (solid line) and compared to an exact calculation (dotted lines).
The second example that will be considered is the time-evolution under the hyperbolic potential from Equation (18) shown in Figure 3(b). Once again, the system is initialised in the ground state of the SHO Hamiltonian centred at the origin. Figure 9(a) shows the comparison between the time-evolution of the wavefunction simulated by the quantum device and the classical device, with a Trotter time-step of $\delta t=0.1$ and a Fock truncation of $n_{\rm max}=60$ on the quantum simulation. To achieve the simulation up to a truncation of 60 the Ket command has been used. The quantum circuit performs well, with the KL divergences showing good agreement for truncations greater than 35, as shown in Figure 8(b).
To fully test the performance of the proposed quantum algorithm, the full circuit was constructed for the potential from Equation (18). Due to the memory constraints on the simulation, this circuit was run using the three-qumode evolver-state preparation circuit from Figure 6 with a Fock truncation of 25. Figure 9(b) shows the time-evolution simulated by the full circuit compared to the classical simulation for a Trotter time-step of $\delta t=0.1$. Good agreement is achieved between the quantum and classical simulations, with the KL divergence of the full circuit matching exactly with the compact simulation using the Ket command, as shown by the exact match between the evolution using the full circuit and the Ket operation in Figure 11. This agreement therefore validates the performance of the method.
@@ 5 Towards quantum field theory
Figure 9: Kullback-Leibler (KL) divergence between the quantum simulation and exact calculation for different times and Fock truncations. The KL divergence quantifies the disparity between probability distributions as a relative entropy (which broadly speaking encodes the information required to get from one distribution to the other). After a sufficient cutoff, the KL divergence exhibits a monotonic behaviour with time.
Given the ability to perform real-time dynamics on a single quantum mechanical state, continuous variable models of quantum computing open up interesting avenues to explore from the perspective of field theory. Both fundamental and effective quantum fields are of paramount importance in many aspects of physics, in particular in particle physics and the Standard Model. Many phenomena, such as strong coupling effects in gauge theory, quantum tunnelling, phase-transitions and other dynamical processes are very hard to study analytically and quantum computing promises to become an important tool, as proposed in the work of References Jordan:2011ci ; Jordan:2012xnu ; Jordan:2014tma ; Jordan_2018 ; Abel:2020qzm ; Abel:2020ebj ; PhysRevA.105.012412 (see References Klco:2018zqz for a more recent review).
The reason that the continuous variable method of quantum computing is an attractive platform for such studies is that fields can be encoded without need for explicitly digitising the field value itself. One can instead simply use the continuous variables to stand for field values. This will turn out to be a great simplification because it then allows the kinetic terms in the field theory Hamiltonian to be constructed using a much smaller number of simple Gaussian gates.
@@
This Section will demonstrate that this can be implemented by outlining a framework for real scalar field theory in 1+1 dimensions. The time dimension will as for the quantum mechanical system be encapsulated by the Trotterised evolution. This is to be accompanied by a single space dimension which is discretised in $M$ qumodes. The expectation values of the fields $\langle\varphi\rangle$ at each point in space will be encoded in the value of the $\langle\hat{x}\rangle$ value on each qumode. In order to avoid confusion the single physical space-dimension will be denoted $r$, and it will be discretised using a one-dimensional lattice of spacing $a$. Thus the field at the $k$ th space position,
@@
It is worth comparing the scaling of this method with that of a discrete system in terms of the required gate operations. Each evolver-gadget contains 3 gates. In addition there are $M$ of the $C_{Z}$ gates. The ancilla circuit for the evolver-state can be reused so this does not need to be included in the circuit count. Thus in total there are $M$ qumodes and $4M$ gates. In a $d$-dimensional system this scales as $(4M)^{d}$ gates. By contrast suppose the field is encoded in a discrete way, with each field value being encoded by $N$ qubits. To make the kinetic cross-terms, every qubit describing the field at a given space point has to be connected to every qubit of the field at the $2d$ nearest neighbour points. Thus one requires at least $M^{d}\times N^{2d}$ gates, even before the potential has been encoded. As is evident it is the gate-count that gets out-of-hand very quickly. Indeed a 3d lattice that is only 10 points on a side with the field encoded in 10 qubits, which gives only 1/32 accuracy assuming a binary encoding of complex values, requires at least a billion gatesOne might suppose that a momentum basis for the embedding could be beneficial, but then the potential would be even more problematic.. The same system encoded on a photonic device, including the potential, would require only $00040^{3}=64,$ 000 gates.
@@ 6 Conclusion
We focused on Continuous-Variable Quantum Computing (CVQC) and its applications in simulating quantum mechanics and quantum field theory. Our investigation stems from recognising the need to surpass classical computational paradigms to deepen our understanding of fundamental physics through quantum mechanical simulations. Our primary objective was to demonstrate the efficacy of CVQC, leveraging the infinite-dimensional Hilbert space of quantum states, for the accurate simulation of quantum mechanics. We achieved this by meticulously constructing a framework to simulate the time evolution of quantum states under arbitrary Hamiltonians using photonic devices. This involved a detailed exploration of Gaussian and non-Gaussian gate operations essential for the manipulation of quantum states encoded in the continuous observables of photons. A pivotal technical achievement in our work is the development of the evolver-state, a specially prepared quantum state that facilitates the desired Trotterised time-evolution of a quantum-mechanical wavefunction. This approach allowed us to simulate the time evolution of quantum systems under arbitrary potentials, using a combination of quantum gate operations and the strategic manipulation of the evolver-state.
@@
The proposed algorithm for simulating the time-evolution of a quantum-mechanical system under the influence of an arbitrary Hamiltonian has been validated against an exact, classical simulation. The circuit shows good agreement with the classical approach for two scenarios, and has been shown to work in full up to a Fock truncation of $n_{\rm max}=25$, limited by memory constraints when simulating the quantum device. In a practical application on a real CVQC device, these limitations would not be present and the full circuit could be achieved. The promising agreement between these simulations underscores the potential of our approach in simulating complex quantum systems.

### v2

@@ 2.2 Creating non-Gaussian operations
Figure 3: The potentials of focus in this study: namely the (a) quartic, (b) $\cosh$ potential and (c) ‘$w$’ potential. In all cases the wave-function is initialised as a Gaussian state centred on the origin.
The circuit parameters are then trained by a classical machine learning routine to fit the output of the circuit to a target state vector. At each step of the training, the loss is calculated by truncating the Fock state expansion from Equation (5). The loss has the form
@@ 3 Implementing time-evolution of wavefunctions
Having collected the required ingredients, we now present a Continuous-Variable Quantum Computing (CVQC) framework for simulating the time-evolution of quantum states under the influence of an arbitrary potential. Section 3.1 describes the principle whereby an arbitrary state can be made to evolve under the influence of an arbitrary Hamiltonian, by using an ancilla qumode initialised with a prescribed non-Gaussian state called the evolver-state. Then Section 3.2 discusses how the evolver state may itself be constructed using the measurement-based approach described in Section 2.2.
@@ 3.1 Schrödinger evolving wavefunctions with arbitrary Hamiltonians
The goal is to be able to evolve any arbitrary input state $|\psi_{\rm in}\rangle$ under the influence of the Hamiltonian, ${\mathcal{H}}$, for which one must devise a photonic circuit that will implement the Schrödinger evolution
@@
The techniques that will be developed here to do this are applicable to any potential, $V$, but in order to have a specific system in mind it is useful to focus on three specific cases. The first is the system with the potential
@@
The depth of the true vacuum of this potential is $V_{\rm min}\approx\varepsilon$, at leading order. The potential is shown in Figure 3(a). Around the origin this potential approximates the simple harmonic oscillator, thus the Gaussian SHO ground state at $x=0$ is an approximate energy eigenstate which will partially decay by tunnelling through the barrier.
@@
This potential is of interest because its expansion around $x=1$ is $V(x)=(x-1)^{2}/2+\ldots$, so to quadratic order it is also the SHO potential. Therefore any deviation from SHO behaviour is directly attributable to the higher-order terms, and moreover this deviation comes from a potential that is not polynomial. The final potential we will consider is
$V(x)~{}=~{}\frac{1}{2}\left(\Theta(1-x)\times x^{2}+\Theta(x-1)\times(x-2)^{2}\right)~{},$ (20)
where $\Theta(x)$ is the Heaviside function, which is shown in Figure 3(c). This ‘$w$’ potential is interesting for a number of reasons. First is the fact that the barrier in the potential is not differentiable, so there is not even a good polynomial approximation for it. The second reason this potential is interesting emerges upon considering solutions to the Schrödinger equation in a potential where the second minimum is missing and instead replaced with a flat region extending from the peak, i.e.  $V(x)~{}=~{}\frac{1}{2}\left(\Theta(1-x)\times x^{2}+\Theta(x-1)\times 1\right)$. In this potential the wavefunction is a bound state and none of it can escape to the right. In other words the ground-state energy of the SHO is smaller than the height of the barrier in the ‘$w$’ potential. Therefore any barrier penetration in the full ‘$w$’ potential of Eq. (20) is entirely due to quantum tunnelling, and the behaviour that will be recovered is characteristically “quantum”.
To implement the Schrödinger evolution of the wavefunction in Equation (17), there are two approximations that will be made. The first is to approximate the Fock expansion from Equation (5). That is
@@
where $A_{n}$ is the coefficient of the Fock state $|n\rangle$, and $n_{\textrm{max}}$ is the Fock state truncation. Thus in principle it is possible to determine the evolution in terms of a matrix acting on the Fock state coefficients, $A_{n}$, by expanding $\hat{x}$ and $\hat{p}$ in terms of the creation and annihilation operators as in Equation (3). Although such Fock truncation is not strictly speaking required for a particular photonic circuit to work, it is required to determine the parameters of the circuit itself, as will become clear.
@@
where the dots denote higher-order commutators. Assuming that $[\hat{p}^{2},H_{1}(\hat{x})]\sim 1$, for such a trotterised evolution one therefore finds that an error accumulates in the exponent of order $N\delta t^{2}$ and hence one requires $\delta t\ll 1/t$ in natural units for the evolution to be accurate. Note that the Trotterisation error is retained as a product of unitary operators, so that it is also unitary***It could in principle be improved by subtracting the leading $\delta t^{2}$ error with more compound Trotter steps, however this will not be necessary..
@@
According to Equations (8) and (14), performing the various manipulations corresponding to these gates and then performing the homodyne measurement $y=0$ with the choice of parameters in Equation (29) yields an evolved state on the $\left|\psi\right\rangle$ qumode of the form
@@
All that remains is to choose the optimal form of the resource function $\left\langle y|\phi_{0}\right\rangle$ in order to maximally suppress the effect of the noise function. There are two possibilities: one can choose a small value of $q$ which ‘freezes’ the value of the function $\left\langle xq|\phi_{0}\right\rangle$, and/or choose a flat resource function. It is convenient to adopt the top-hat function as the idealised resource function:
@@
$|\psi\mbox{\small$(\delta t)$}\rangle~{}=~{}\frac{\Theta(L/2-q|\hat{x}|)}{L}~{}e^{-i{\cal H}(\hat{x})\,\delta t}|\psi_{\rm in}\rangle~{}.$ (32)
Note that smaller values of $q$ allow larger domains in $x$ of valid evolution.
@@ 3.2 Learning the evolver-state
To fully implement the procedure outlined in Section 3.1 on a photonic device, the non-Gaussian evolver-state, $\left|\phi\right\rangle$, must be prepared. Here, the measurement-based approach from Section 2.2 will be used, in which the circuit parameters of a circuit such as that in Figure 5 are trained against the target state from Equation (34). To achieve a good fit to the target, a circuit constructed from three-qumodes and 10 iterations of the layer method will be used. At each layer, the displaced-squeezed states pass through an interferometer which entangles the system. Following in the rectangular architecture from Reference Clements2016 only three beamsplitters are required in each layer for the three-qumode case. Each gate operation has two parameters, thus the full circuit has 180 trainable variables. The training of the gate parameters has been restricted to values which are experimentally realisable PhysRevLett.117.110801 . The values that the PNR measurements are post-selected on have not been included as trainable parameters and have instead been chosen to be $m_{0}=m_{1}=5$. Reference PhysRevA.100.052301 makes a detailed investigation into maximising success when creating non-Gaussian states using measurement-based quantum computing approaches. A schematic of the circuit diagram used to create the desired resource state for the time-evolution of an arbitrary Hamiltonian is shown in Figure 6.
@@
4 Results: Quantum mechanics on photonics versus numerically evaluated quantum mechanics
@@
Figure 8: The time-evolution of a quantum system with the asymmetric quartic-potential of Equation (18) with (a) $\varepsilon=0.1$ and (b) $\varepsilon=0.5$, generated by the photonic quantum simulator with a Fock truncation of 60 (solid line) and compared to an exact calculation (dotted lines).
The second example that we consider is the time-evolution under the hyperbolic potential from Equation (19) shown in Figure 3(b). Once again, the system is initialised in the ground state of the SHO Hamiltonian centred at the origin. Figure 10(a) shows the comparison between the time-evolution of the wavefunction simulated by the quantum device and the classical device, with a Trotter time-step of $\delta t=0.1$ and a Fock truncation of $n_{\rm max}=60$ on the quantum simulation. To achieve the simulation up to a truncation of 60 the Ket command has been used. The quantum circuit performs well, with the KL divergences showing good agreement for truncations greater than 35, as shown in Figure 9(b).
Finally we consider evolution in the ‘$w$’ potential of Eq. (20) shown in Figure 3(c). Again, the system is initialised in the ground state of the SHO Hamiltonian centred at the origin. Figure 11 shows the comparison between the time-evolution of the wavefunction simulated by the quantum device and the classical device, with a Trotter time-step of $\delta t=0.1$ and a Fock truncation of $n_{\rm max}=60$ on the quantum simulation. The evolution is extraordinarily accurate with this potential. Indeed the KL divergences, which are shown in Figure 9(c), are extremely small for a sufficiently large Fock truncation.
It is interesting to ask why the evolution in the ‘$w$’ potential should be so much more accurate. Recall that the evolution in this case is expected to initially be dominated by tunnelling, implying that the penetration of the barrier (of height $V(1)=0.5$) is driven by exponential tails of the wavefunction. This in turn implies that the whole wavefunction is a bound state of the double well that must be exponentially suppressed beyond $x<-1$ and $x>3$. It is therefore insensitive to the edges of the top-hat resource state (which are well outside this range) and the noise-factor $\langle qx|\phi_{0}\rangle$ appearing in Eq. (31), in contrast with the situation in the other two potentials.
To fully test the performance of the proposed quantum algorithm, the full circuit was constructed for the cosh potential of Eq. (19). Due to the memory constraints on the simulation, this circuit was run using the three-qumode evolver-state preparation circuit from Figure 6 with a Fock truncation of 25. Figure 10(b) shows the time-evolution simulated by the full circuit compared to the classical simulation for a Trotter time-step of $\delta t=0.1$. Good agreement is achieved between the quantum and classical simulations, with the KL divergence of the full circuit matching exactly with the compact simulation using the Ket command, as shown by the exact match between the evolution using the full circuit and the Ket operation in Figure 12. This agreement therefore validates the performance of the method.
@@ 5 Towards quantum field theory
Figure 9: The Kullback-Leibler (KL) divergence between the quantum simulation and exact calculation for different evolution times and Fock truncations for the (a) quartic, (b) $\cosh$ and (c) ‘$w$’ potentials. The KL divergence quantifies the disparity between probability distributions as a relative entropy (which broadly speaking encodes the information required to get from one distribution to the other). After a sufficient cutoff, the KL divergence exhibits a monotonic behaviour with time.
Given the ability to perform real-time dynamics on a single quantum-mechanical state, continuous-variable models of quantum computing open up interesting avenues to explore from the perspective of field theory. Both fundamental and effective quantum fields are of paramount importance in many aspects of physics, in particular in particle physics and the Standard Model. Many phenomena, such as strong coupling effects in gauge theory, quantum tunnelling, phase-transitions and other dynamical processes are very hard to study analytically and quantum computing promises to become an important tool, as proposed in the work of References Jordan:2011ci ; Jordan:2012xnu ; Jordan:2014tma ; Jordan_2018 ; Abel:2020qzm ; Abel:2020ebj ; PhysRevA.105.012412 (see References Klco:2018zqz for a more recent review).
The reason that the continuous-variable method of quantum computing is an attractive platform for such studies is that fields can be encoded without need for explicitly digitising the field value itself. One can instead simply use the continuous variables to stand for field values. This will turn out to be a great simplification because it then allows the kinetic terms in the field theory Hamiltonian to be constructed using a much smaller number of simple Gaussian gates.
@@
This Section will demonstrate that this can be implemented by outlining a framework for real scalar field theory in 1+1 dimensions. The time dimension will as for the quantum-mechanical system be encapsulated by the Trotterised evolution. This is to be accompanied by a single space dimension which is discretised in $M$ qumodes. The expectation values of the fields $\langle\varphi\rangle$ at each point in space will be encoded in the value of the $\langle\hat{x}\rangle$ value on each qumode. In order to avoid confusion the single physical space-dimension will be denoted $r$, and it will be discretised using a one-dimensional lattice of spacing $a$. Thus the field at the $k$ th space position,
@@
Figure 11: The time-evolution of a quantum system under the influence of the ‘$w$’ potential of Equation (20) generated by the photonic quantum simulator (solid lines) compared to an exact calculation (dotted lines).
@@
It is worth comparing the scaling of this method with that of a discrete system in terms of the required gate operations. Each evolver-gadget contains 3 gates. In addition there are $M$ of the $C_{Z}$ gates. The ancilla circuit for the evolver-state can be reused so this does not need to be included in the circuit count. Thus in total there are $M$ qumodes and $4M$ gates. In a $d$-dimensional system this scales as $(4M)^{d}$ gates. By contrast suppose the field is encoded in a discrete way, with each field value being encoded by $N$ qubits. To make the kinetic cross-terms, every qubit describing the field at a given space point has to be connected to every qubit of the field at the two-dimensional nearest-neighbour points. Thus one requires at least $M^{d}\times N^{2d}$ gates, even before the potential has been encoded. As is evident it is the gate-count that gets out-of-hand very quickly. Indeed a three-dimensional lattice that is only 10 points on a side with the field encoded in 10 qubits, which gives only 1/32 accuracy assuming a binary encoding of complex values, requires at least a billion gatesOne might suppose that a momentum basis for the embedding could be beneficial, but then the potential would be even more problematic.. The same system encoded on a photonic device, including the potential, would require only $00040^{3}=64,$ 000 gates.
@@ 6 Conclusion
We focused on Continuous-Variable Quantum Computing (CVQC) and its applications in simulating quantum mechanics and quantum field theory. Our investigation stems from recognising the need to surpass classical computational paradigms to deepen our understanding of fundamental physics through quantum-mechanical simulations. Our primary objective was to demonstrate the efficacy of CVQC, leveraging the infinite-dimensional Hilbert space of quantum states, for the accurate simulation of quantum mechanics. We achieved this by meticulously constructing a framework to simulate the time evolution of quantum states under arbitrary Hamiltonians using photonic devices. This involved a detailed exploration of Gaussian and non-Gaussian gate operations essential for the manipulation of quantum states encoded in the continuous observables of photons. A pivotal technical achievement in our paper is the development of the evolver-state, a specially prepared quantum state that facilitates the desired Trotterised time-evolution of a quantum-mechanical wavefunction. This approach allowed us to simulate the time evolution of quantum systems under arbitrary potentials, using a combination of quantum gate operations and the strategic manipulation of the evolver-state.
@@
The proposed algorithm for simulating the time-evolution of a quantum-mechanical system under the influence of an arbitrary Hamiltonian has been validated against an exact, classical simulation. The circuit shows good agreement with the classical approach for three scenarios, and has been shown to work in full up to a Fock truncation of $n_{\rm max}=25$, limited by memory constraints when simulating the quantum device. In a practical application on a real CVQC device, these limitations would not be present and the full circuit could be achieved. The promising agreement between these simulations underscores the potential of our approach in simulating complex quantum systems.